james savings account earns 7% interest compounded quarterly. james deposited $950. find the balance after 22 months
Found 2 solutions by stanbon, robertb:
Answer by stanbon(75887)

james savings account earns 7% interest compounded quarterly. james deposited $950. find the balance after 22 months
-----
A(t) = P(1+(r/n))^(nt)
----------
A(22/12) = 950(1+(0.07/4))^(4*(22/12))
----------
A(22/12) = 950(1.1357)
A(22/12) = $1078.88
========================
Cheers,
Stan H.

Answer by robertb(5830)

P = 950, r = 0.07, n = 4, t = 22/12 = 11/6. Then